# Constrained Decoding for Robotics Foundation Models

### v3

@@ Abstract
Recent advances in the development of robotic foundation models have led to promising end-to-end and general-purpose capabilities in robotic systems. Trained on vast datasets of simulated and real-world trajectories, these models map multimodal observations directly to action sequences for physical execution. Despite promising real-world capabilities, these models are still data-driven and, therefore, lack explicit notions of behavioral correctness. We address this gap by introducing SafeDec, a constrained decoding framework for autoregressive, robot foundation models that enforces invariant safety specifications on candidate action trajectories. Task-specific safety rules are expressed as Signal Temporal Logic (STL) formulas and are enforced at inference time with minimal overhead. Our method ensures that generated actions provably satisfy STL specifications under assumed dynamics at runtime without retraining , while remaining agnostic of the underlying policy. We evaluate SafeDec on tasks from the CHORES benchmark for state-of-the-art generalist policies (e.g., SPOC, Flare, PoliFormer) across hundreds of procedurally generated environments and show that our decoding-time interventions are useful not only for filtering unsafe actions but also for conditional action generation. Videos are available at constrained-robot-fms.github.io.
@@ 1 Introduction
Recent advances in developing Robotics Foundation Models (RFM) have enabled general-purpose policies that map multi-modal inputs such as RGB images, natural language instructions, and proprioceptive inputs to action sequences Hu et al. [1]. Shortest Path Oracle Clone (SPOC) [2], PoliFormer [3], Flare [4] and OpenVLA [5] exhibit impressive generalization in navigation and manipulation tasks and serve as versatile robot controllers for real-world deployment contexts. However, these models are primarily data-driven and lack any explicit notion of safety. Although these models may implicitly exhibit safety-related behaviors depending on the patterns in their training data, there is no formal guarantee that models will consistently behave safely in all situations. This serves as a limiting factor for deploying these foundation models in the physical world where rule compliance and regulatory safety rule adherence are crucial.
@@
Formal specifications have long been used to specify safety requirements for robotic deployments [6, 7]. Temporal logics (TL) [8]  can capture safety constraints on robot behavior, such as “remain within the permitted region zones and avoid dangerous obstacles”. Although TL has seen success in classical robotic planning for safety constraint satisfaction, its use for enforcing safety for large transformer-based robot policies remains limited. Additionally, retraining or fine-tuning these large pre-trained models to directly embed temporal logic specification is challenging [9]. First, retraining models is a costly endeavor in terms of computational resources and data requirements. Moreover, due to the stochastic nature of these models, it is difficult to guarantee strict satisfaction of safety constraints through training alone. Hence, there is a pressing need for methods that can enforce safety specifications efficiently at inference time without disrupting the model’s pre-trained behavior.
@@
In the field of natural language processing, syntactic constraints have been successfully enforced by applying constrained decoding at inference time [10, 11, 12]. These approaches typically mask out tokens that violate a syntactic constraint defined over token sequences. For example, regular expressions (regex) represent a widely used form of syntactic constraint, requiring that generated token sequences conform to predefined structural patterns [10, 11]. Inspired by this line of work, we extend the paradigm of constrained decoding to enforce safety constraints over action trajectories in dynamical systems and propose safety specification aligned decoding (SafeDec) for transformer based policies that ensures generated action sequences provably satisfy Signal Temporal Logic (STL) [13] specifications under assumed dynamics. Our key insight is that decoding-time interventions can be used not just to filter unsafe actions, but to condition the generation process itself on specification satisfaction. This conditioning is critical because it steers the model toward generating safety specification satisfying actions rather than relying on post hoc rejection. SafeDec reduces risk of infeasible outputs while preserving the original action distribution of the model. To enforce such specifications, we leverage the formal semantics of STL to evaluate candidate actions at runtime and mask those that lead to future violations. Our method is agnostic to the underlying foundation model, requiring only two properties: (1) access to the decoding-layer logits during inference, and (2) access to an approximate dynamics model to predict future states. To efficiently evaluate STL specifications at inference time, we use a high-performance computational graph based library STLCG++ [14]. To the best of our knowledge, this is the first work to enforce formal safety on transformer-based robotic policies at inference time using constrained decoding.
Our main contributions are as follows: First, we introduce the novel problem of constrained decoding for transformer-based policies under safety STL specifications (Section 3.1). Second, we propose an inference-time technique that reweights or masks candidate actions using STL satisfaction scores in (Section 3.2 and 3.3). Finally, we demonstrate the effectiveness of our method on state-of-the-art object navigation models without modifying model parameters (Section 4).
@@ 2.1 Signal Temporal Logic
Signal Temporal Logic (STL) is an expressive framework for defining properties and reasoning over continuous time real valued signals [15]. Formally, $(\mathbf{s},t)\models\phi$ denotes that a signal $\mathbf{s}$ satisfies the STL formula $\phi$ at time $t$. An atomic predicate of an STL formula is represented by inequalities of the form $\mu(\mathbf{s}(t))>0$. The truth value of the predicate $\mu$ is equivalent to $\mu(\mathbf{s}(t))>0$. Note that with slight abuse of notation, $\mu$ represents both the predicate and a function of the trajectory $\mathbf{s}(t)$. Any STL formula consists of Boolean and temporal operations on these predicates, and the syntax of STL formulas is defined recursively as follows:
@@
Apart from the Boolean semantics, quantitative semantics are defined for a signal to compute a real-valued metric indicating robustness, i.e., the strength of satisfaction or violation. For the sake of brevity, the definition of robustness is provided in Appendix A.1.
@@ 2.2 Constrained Decoding in Transformers
A large variety of autoregressive transformer-based models generate final outputs by producing a probability distribution over the model vocabulary at each timestep. This distribution is generated by performing a softmax operation over the model’s last hidden layer. Then, through the process of decoding, tokens are selected to maximize the overall likelihood of an output sequence. In standard decoding, this maximization can be performed by either greedily selecting the most probable token at each step or by using a beam search to maintain multiple high-likelihood candidates. However, this often leads to degenerate output sequences that are repetitive [16]. A common approach is to use sampling strategies like top $k$ [17], and nucleus sampling [16] that introduce stochasticity to encourage more diverse outputs. Constrained decoding [18] modifies this probabilistic selection by pruning invalid tokens to ensure that the generated sequences satisfy predefined constraints. These constraints are often syntactic, such as regular expressions, JSON formatting, or programming language grammars [19]. There is also recent work on enforcing semantic constraints that ensure coherence of the output or alignment with specific knowledge bases [20]. Formally, constrained decoding can be seen as maximizing the probability of the output sequence subject to a constraint $\mathcal{C}$: $\arg\max_{y\in\mathcal{Y}_{\mathcal{C}}}P(y\mid x)$ where $\mathcal{Y}_{\mathcal{C}}$ is the set of sequences satisfying $\mathcal{C}$.
@@ 3 Specification-Guided Constrained Decoding
Figure 1: Overview of our specification aligned decoding framework. Given multimodal inputs (e.g., RGB images and natural language instructions), a pretrained robot foundation model (e.g., SPOC) generates candidate actions. These actions are filtered or reweighted by the constrained decoding technique based on STL safety constraints. In the figure, green markers denote target object locations, while red zones represent regions to avoid.
In this section, we introduce a novel problem formulation for SafeDec in RFMs. First, we highlight the challenge in specification checking for these models in contrast to traditional syntactical constraint checking adopted by LLMs, and our solution to remedy it. Then, we propose two novel inference-time techniques for specification aligned decoding: Hard Constrained Decoding (HCD) and Robustness Constrained Decoding (RCD).
@@ 3.1 Problem Statement
In contrast, RFMs operate in a physical environment and constraints (captured via temporal logic) are defined over state variables. Since a large class of end-to-end policies solely propose action sequences, specification checking can only be performed as actions are executed and the environment is simulated forward. In this case, constraint checking cannot be done solely in the token space and requires environmental feedback or a dynamics stepping function. Hence, we leverage an approximate first order dynamics function to compute specification satisfaction of different action sequences proposed by these policies.
@@
In this work, we focus on RFMs that generate actions based on multi-modal inputs, including sensor observations (e.g., RGB, depth, LiDAR) and natural language instructions. Let $\mathcal{I}_{t}$ represent the aggregated input at timestep $t$. Given the history of encoded inputs up to time $t$, a transformer-based policy parameterized by $\theta$ predicts embeddings for the next TtT{-}t actions:
@@
Most of the existing techniques for specification enforcement perform posthoc manipulation of proposed actions through filtering or rejecting action sequences that violate the specification $\varphi$. Although manipulation after sampling can ensure specification satisfaction, it can lead to distorting the model’s learned distribution, producing low-likelihood output. This undermines the inductive biases learned during pretraining and leads to degenerate, brittle behaviors. A similar problem was highlighted when ensuring compliance with logical constraints for large language models (LLM) in Park et al. [21]. Additionally, these models decode actions sequentially, where each action $a_{t}$ is conditioned on previously generated tokens $a_{<t}$. Posthoc manipulation can disrupt this causal chain and lead to a mismatch between the model’s internal hidden state and the executed sequence. Hence, we propose the following problem statement:
@@
In this work, we propose a technique to overcome the drawbacks of post-hoc safety enforcement methods (such as filtering) by leveraging constrained decoding techniques. Specifically, we propose SafeDec : A constrained decoding strategy that integrates STL specifications into the foundation model action selection process itself, ensuring satisfaction without distorting the model’s distribution.
@@ 3.2 Hard Constrained Decoding
As highlighted in the background section, in the final layer, predictions are detokenized and a projection layer converts the embeddings into logits over the vocabulary space. These logits are further converted into a probability distribution using a softmax operation. In prior work, for structured output generation in LLMs, some invalid tokens are masked based on syntactical constraints or other criteria [19, 21]. This is done by setting their logit value as -$\infty$ before the softmax operation is applied. For HCD, we use a similar approach as constrained decoding literature [19] and mask out predicted action tokens that violate our given STL specification $\varphi$ during sequential generation. Formally, to enforce the STL specification $\varphi$ during sequential generation, we adjust the logits at each timestep $t+k$ as follows:
@@
$z_{t+k}^{(i)}=\begin{cases}-\infty,&\text{if }\hat{x}_{t+k}^{(i)}=f(x_{t+k-1},\hat{a}_{t+k}^{(i)})\text{ violates }\varphi\\ z_{t+k}^{(i)},&\text{otherwise}\end{cases}$
Here $t$ is the current decision step, $k$ is an index for the look-ahead step $t+k$ within a planning horizon of length $T$ ($k\in[1..T]$), $\hat{a}_{t+k}^{(i)}$ is the action mapping to the token i and $\hat{x}_{t+k}^{(i)}$ is the next state value upon taking this action. This next state is elicited using a simple dynamics model ($f$) as highlighted in the previous section. Adjusting logits in this fashion ensures that any invalid token with respect to the safety specification will have zero probability of being selected after applying the softmax function.
@@ 3.3 Robustness Constrained Decoding
HCD ensures compliance but can lead to compromising task success, which can be undesirable. A similar tradeoff was observed by Liu et al. [22] when probability space-steering preserved model fluency while reducing toxic continuations compared with hard-filtering strategies that inflated perplexity and eroded diversity. Hence, we propose an alternative approach, called RCD, where we leverage the quantitative semantics of STL specifications (robustness). Unlike HCD, which applies hard masking to completely remove unsafe actions, RCD softly guides the model toward safer actions by incorporating robustness scores that reflect the degree of satisfaction of $\varphi$. This is similar to the approach proposed in Liu et al. [22] where the next-token distributions were re-weighted based on the utility scores provided by another language model. Our utility scores are quantified by the robustness function( $\rho(x_{t},\varphi)$ ) that returns a real-valued score indicating how well a predicted state satisfies the specification. Positive robustness values denote specification satisfaction, while negative values capture the degree of violation. A formal definition of robustness in line with the STL quantitative semantics is provided in Appendix A.1.
First, we compute a robustness score for each candidate action: $r_{t+k}^{(i)}=\rho(\hat{x}_{t+k}^{(i)},\varphi)$ where $\rho(\cdot,\varphi)$ is the STL robustness metric, and $\hat{x}_{t+k}^{(i)}$ is the predicted next state under action $\hat{a}_{t+k}^{(i)}$. This robustness score $r_{t+k}^{(i)}$ quantifies how well each candidate action satisfies the specification $\varphi$. These scores are then converted into weights using exponential scaling: $w_{t+k,i}=\exp(\alpha\cdot r_{t+k,i})$ where $\alpha$ is a temperature parameter that adjusts the sharpness of the bias. We use these weights to shift the original logits: $\tilde{z}_{t+k,i}=z_{t+k,i}+\beta\cdot w_{t+k,i}$ where $\beta$ is a hyperparameter that modulates the trade-off between specification adherence and the original task objective. Finally, we obtain the action distribution by applying softmax over the adjusted logits: $p_{t+k}=\text{softmax}(\tilde{\mathbf{z}}_{t+k})$

### v4

# SafeDec: Constrained Decoding for Safe Autoregressive Generalist Robot Navigation Policies

@@ Abstract
Recent advances in end-to-end, multi-task robot policies based on transformer models have demonstrated impressive generalization to real-world embodied navigation tasks. Trained on vast datasets of simulated and real-world trajectories, these policies map multimodal observations directly to action sequences for physical execution. Despite promising real-world capabilities, these models are still data-driven and, therefore, lack explicit notions of behavioral correctness. We address this gap by introducing SafeDec, a constrained decoding framework for autoregressive, transformer-based robot navigation policies that enforces safety specifications expressed as Signal Temporal Logic (STL) formulas. Our method ensures that generated actions provably satisfy STL specifications under assumed dynamics at runtime without retraining while remaining agnostic of the underlying policy. We evaluate SafeDec on tasks from the CHORES benchmark for state-of-the-art embodied navigation policies across hundreds of procedurally generated environments and show that our decoding-time interventions are useful not only for filtering unsafe actions but also for conditional action generation. Videos are available at constrained-robot-fms.github.io.
Machine Learning, Safe Robotics, Constrained Decoding, Robot Navigation
@@ 1 Introduction
Recent advances in developing large transformer-based models for embodied navigation have enabled general-purpose policies that map multi-modal inputs such as RGB images, natural language instructions, and proprioceptive inputs to action sequences (Hu et al., 2023). For example, Shortest Path Oracle Clone (SPOC) (Ehsani et al., 2024), PoliFormer (Zeng et al., 2025), Flare (Hu et al., 2025), and LEO (Huang et al., 2024) exhibit impressive zero-shot generalization for embodied navigation tasks such as open vocabulary object-goal navigation (“find a mug”) and attribute-conditioned navigation (“locate the chair closest to the refrigerator in the kitchen”). These models have also shown promising transfer for robot navigation in real-world settings (Zeng et al., 2025; Huang et al., 2024). However, these policies are primarily data-driven and lack any explicit notion of safety. Although policies may implicitly exhibit safety-related behaviors depending on patterns in their training data, there is no formal guarantee that they will consistently behave safely in all situations. This limitation hinders the deployment of such policies in the physical world where adherence to explicit safety rules and regulatory constraints is essential.
Formal specifications have long been used to specify contextual safety requirements for robotic deployments (Menghi et al., 2019; Farrell et al., 2018). Specifically, temporal logics (TL) (Pnueli, 1977)  provide a structured language for specifying rich safety constraints on robot behavior, ranging from simple invariants (e.g., remaining within permitted operational zones or avoiding dangerous obstacles) to contextual and history-dependent specifications. For example, a robot may be required to avoid certain regions until visiting another region (temporal ordering), permanently avoid a region after entering a restricted area (conditional safety) or never return to a region once it has exited. Such specifications naturally capture real-world operational rules that depend not only on the robot’s current state, but also on what situations the robot has previously encountered.
@@
Although TL has seen success in classical robotic planning and reinforcement learning for safety constraint satisfaction, its use for enforcing safety for large transformer-based robot policies remains limited (Manganaris et al., 2026). Additionally, retraining or fine-tuning these large pre-trained policies to directly embed such non-Markovian TL safety specifications is challenging (Kapoor et al., 2024). First, retraining policies is a costly endeavor in terms of computational resources and data requirements. Moreover, due to the stochastic nature of these models, it is difficult to guarantee strict satisfaction of safety constraints through training alone. Hence, there is a pressing need for methods that can reliably enforce safety specifications at inference time without policy retraining.
@@
In the field of natural language processing, syntactic constraints have been successfully enforced by applying constrained decoding at inference time (Willard and Louf, 2023; Beurer-Kellner et al., 2023; AI, 2023). These approaches typically mask out tokens that violate a syntactic constraint defined over token sequences. For example, regular expressions (regex) represent a widely used form of syntactic constraint, requiring that generated token sequences conform to predefined structural patterns (Willard and Louf, 2023; Beurer-Kellner et al., 2023). Inspired by this line of work, we extend the paradigm of constrained decoding to enforce contextual safety constraints over action trajectories in dynamical systems. We propose safety specification aligned decoding (SafeDec) for transformer based policies that ensures generated action sequences provably satisfy Signal Temporal Logic (STL) (Maler and Nickovic, 2004) specifications under assumed dynamics. Our key insight is that decoding-time interventions can be used not just to filter unsafe actions, but to condition the generation process itself on specification satisfaction. This conditioning is critical because it steers the model toward generating safety specification satisfying actions rather than relying on post hoc rejection. To enforce safety specifications, we leverage the formal semantics of STL to evaluate candidate actions at runtime and mask those that lead to future violations. Our method is agnostic to the underlying foundation model, requiring only two properties: (1) access to the decoding-layer logits during inference, and (2) access to an approximate dynamics model to predict future states. In this work, we focus on safety enforcement for embodied navigation policies in indoor environments. We also focus on STL safety (Alpern and Schneider, 1987) specifications without liveness operators, since such specifications are sufficient to model a wide range of safety constraints encountered in robot learning (He et al., 2024; Yun et al., 2025; Kapoor et al., 2025b; Zhao et al., 2024). To the best of our knowledge, this is the first work to enforce STL safety specifications on transformer-based navigation policies at inference time using constrained decoding.
@@ 2.1 Signal Temporal Logic
Signal Temporal Logic (STL) is an expressive framework for defining properties and reasoning over continuous time real-valued (Maler and Nickovic, 2004) . Formally, $(\mathbf{s},t)\models\phi$ denotes that a signal $\mathbf{s}$ satisfies the STL formula $\phi$ at time $t$. An atomic predicate of an STL formula is represented by inequalities of the form $\mu(\mathbf{s}(t))>0$. The truth value of the predicate $\mu$ is equivalent to $\mu(\mathbf{s}(t))>0$. Note that with slight abuse of notation, $\mu$ represents both the predicate and a function of the trajectory $\mathbf{s}(t)$. Any STL formula consists of Boolean and temporal operations on predicates, and the syntax of STL formulas is defined recursively as follows:
@@
Apart from Boolean semantics, quantitative semantics are defined for a signal to compute a real-valued metric called robustness, i.e., the strength of satisfaction or violation. The definition of robustness is provided in Appendix A.1.
@@ 2.2 Constrained Decoding in Transformers
A large variety of autoregressive transformer-based models generate final outputs by producing a probability distribution over the model vocabulary at each timestep. Then, through the process of decoding, tokens are selected to maximize the overall likelihood of an output sequence. In standard decoding, this maximization can be performed by either greedily selecting the most probable token at each step or by using a beam search to maintain multiple high-likelihood candidates. However, this often leads to degenerate output sequences that are repetitive (Holtzman et al., 2019). A common approach is to use sampling strategies like top $k$ (Fan et al., 2018), and nucleus sampling (Holtzman et al., 2019) that introduce stochasticity to encourage more diverse outputs. Constrained decoding (Hokamp and Liu, 2017) modifies this probabilistic selection by pruning invalid tokens to ensure that the generated sequences satisfy predefined constraints. These constraints are often syntactic, such as regular expressions, JSON formatting, or programming language grammars (Welleck et al., 2024). There is also recent work on enforcing semantic constraints that ensure coherence of the output or alignment with specific knowledge bases (Peyrard et al., 2024). Formally, constrained decoding can be seen as maximizing the probability of the output sequence subject to a constraint $\mathcal{C}$: $\arg\max_{y\in\mathcal{Y}_{\mathcal{C}}}P(y\mid x)$ where $\mathcal{Y}_{\mathcal{C}}$ is the set of sequences satisfying $\mathcal{C}$.
@@ 3 Specification-Guided Constrained Decoding
Figure 1: Overview of our specification aligned decoding framework. Given multimodal observations (RGB images and language goals), a pretrained transformer-based navigation policy generates candidate actions. Our constrained decoder then filters or reweights these actions using robustness scores from a user-defined STL safety specification. Left: Original action probabilities (red) from the pretrained policy are modified by SafeDec using STL robustness: actions predicted to violate the safety specification receive reduced weights, while safer actions are boosted. The chosen safe action and adjusted probabilities (green) are highlighted. Right: An example navigation episode where the robot starts in the kitchen (white dot) and moves toward the bedroom to locate the target object (green marker middle-right), while avoiding user-defined hazardous regions (red zones) throughout the trajectory.
In this section, we introduce a novel problem formulation for SafeDec in autoregressive transformer-based policies. First, we highlight the challenge in specification checking for these policies in contrast to traditional syntactical constraint checking adopted by LLMs, and our solution to remedy it. Then, we propose two novel inference-time techniques for specification aligned decoding: Hard Constrained Decoding (HCD) and Robustness Constrained Decoding (RCD).
@@ 3.1 Problem Statement
In contrast, transformer-based policies operate in a physical environment and constraints (captured via temporal logic) are defined over state variables. Since a large class of end-to-end policies solely propose action sequences, specification checking can only be performed as actions are executed and the environment is simulated forward. In this case, constraint checking cannot be done solely in the token space and requires environmental feedback or a dynamics stepping function. Hence, we leverage an approximate first order dynamics function to compute specification satisfaction of different action sequences proposed by these policies.
@@
In this work, we focus on robot policies that generate actions based on multi-modal inputs, including sensor observations (e.g., RGB, depth, LiDAR) and natural language instructions. Let $\mathcal{I}_{t}$ represent the aggregated input at timestep $t$. Given the history of encoded inputs up to time $t$, a transformer-based policy parameterized by $\theta$ predicts embeddings for the next TtT{-}t actions:
@@
Most of the existing techniques for specification enforcement perform post-hoc manipulation of proposed actions (after sampling from the policy) by filtering or rejecting action sequences that violate the specification $\varphi$. Although manipulation after sampling can ensure specification satisfaction, it can favor safe but low-likelihood actions and unintentionally suppress feasible safe behaviors that the model would have selected if safety constraints had been considered pre-sampling. This undermines the inductive biases learned during pretraining and lead to degenerate and brittle behaviors. A similar problem was highlighted when ensuring compliance with logical constraints for large language models (LLM) in Park et al. (2024). Additionally, these models decode actions sequentially, where each action $a_{t}$ is conditioned on previously generated tokens $a_{<t}$. Post hoc manipulation can disrupt this causal chain and lead to a mismatch between the model’s internal hidden state and the executed sequence. Hence, we propose the following problem statement:
@@
In this work, we propose a technique to overcome the drawbacks of post-hoc safety enforcement methods (such as filtering) by leveraging constrained decoding techniques. Specifically, we propose SafeDec : A constrained decoding strategy that integrates STL specifications into the policy’s action selection process, ensuring satisfaction while remaining as close as possible to the base model distribution.
@@ 3.2 Hard Constrained Decoding
As highlighted in the background section, in the final layer, predictions are detokenized and a projection layer converts the embeddings into logits over the vocabulary space. These logits are further converted into a probability distribution using a softmax operation. In prior work, for structured output generation in LLMs, some invalid tokens are masked based on syntactical constraints or other criteria (Welleck et al., 2024; Park et al., 2024). This is done by setting their logit value as -$\infty$ before the softmax operation is applied. For HCD, we use a similar approach as constrained decoding literature (Welleck et al., 2024) and mask out predicted action tokens that violate our given STL specification $\varphi$ during sequential generation. Formally, to enforce the STL specification $\varphi$ during sequential generation, we adjust the logits at each timestep $t+k$ as follows:
@@
$z_{t+k}^{(i)}=\begin{cases}-\infty,&\text{if }\hat{\tau}^{(i)}_{0:t+k}=(x_{0},x_{1},..,\hat{x}_{t+k}^{(i)})\text{ violates }\varphi\\ z_{t+k}^{(i)},&\text{otherwise}\end{cases}$
where $\hat{x}_{t+k}^{(i)}=f(\hat{x}_{t+k-1},\hat{a}_{t+k}^{(i)})$.
Here, $t$ is the current decision step, $k$ is an index for the look-ahead step $t+k$ within a planning horizon of length $T$ ($k\in[1..T]$), $\hat{a}_{t+k}^{(i)}$ is the action mapping to the token i and $\hat{x}_{t+k}^{(i)}$ is the next state value upon taking this action. This next state is elicited using a simple dynamics model ($f$) as highlighted in the previous section. Adjusting logits in this fashion ensures that any invalid token with respect to the safety specification will have zero probability of being selected after applying the softmax function.
Table 1: STL specifications used in our evaluation. Each region $R_{i}$ denotes a spatial predicate over $(x,z)$.
Spec name STL formula Description
$\phi_{\text{avoid}}$ $\mathbf{G}\!\left(\bigwedge_{i=1}^{N}\neg R_{i}\right)$ Avoidance: the robot must always stay outside all unsafe regions.
$\phi_{\text{ordered}}$ $\mathbf{G}(\neg R_{2})\ \vee\ (\neg R_{2}\ \mathbf{U}\ R_{1})$ Weak-until ordering: the robot must avoid region $R_{2}$ until reaching region $R_{1}$, or avoid $R_{2}$ forever.
$\phi_{\text{no\_return}}$ $\mathbf{G}\big(R_{0}\Rightarrow\mathbf{G}(\neg R_{0}\Rightarrow\mathbf{G}\neg R_{0})\big)$ Conditional no-return: if the robot is in region $R_{0}$, then once it leaves $R_{0}$, it may never re-enter.
$\phi_{\text{geofence}}$ $\mathbf{G}\!\left(\bigvee_{i=1}^{N}R_{i}\right)$ Geofencing: the robot must always remain within at least one allowed region.
@@ 3.3 Robustness Constrained Decoding
HCD ensures compliance but can lead to compromising task success, which can be undesirable. A similar tradeoff was observed by Liu et al. (2021) when probability space-steering preserved model fluency while reducing toxic continuations compared with hard-filtering strategies that inflated perplexity and eroded diversity. Hence, we propose an alternative approach, called RCD, where we leverage the quantitative semantics of STL specifications (robustness). Unlike HCD, which applies hard masking to completely remove unsafe actions, RCD softly guides the model toward safer actions by incorporating robustness scores that reflect the degree of satisfaction of $\varphi$. This is similar to the approach proposed in Liu et al. (2021) where the next-token distributions were re-weighted based on the utility scores provided by another language model. Our utility scores are quantified by the robustness function( $\rho(\langle x_{0},x_{1},\dots,x_{t}\rangle,\varphi)$ ) that returns a real-valued score indicating how well a predicted state satisfies the specification. Positive robustness values denote specification satisfaction, while negative values capture the degree of violation. A formal definition of robustness in line with the STL quantitative semantics is provided in Appendix A.1.
First, we compute a robustness score for each candidate action: $r_{t+k}^{(i)}=\rho(\langle x_{0},x_{1},\dots,x_{t+k-1},\hat{x}_{t+k}^{(i)}\rangle,\varphi)$ where $\rho(\cdot,\varphi)$ is the STL robustness metric, and $\hat{x}_{t+k}^{(i)}$ is the predicted next state under action $\hat{a}_{t+k}^{(i)}$. This robustness score $r_{t+k}^{(i)}$ quantifies how well each candidate action satisfies the specification $\varphi$. These scores are then converted into weights using exponential scaling: $w_{t+k,i}=\exp(\alpha\cdot r_{t+k}^{(i)})$ where $\alpha$ is a temperature parameter that adjusts the sharpness of the bias. We use these weights to shift the original logits: $\tilde{z}_{t+k,i}=z_{t+k,i}+\beta\cdot w_{t+k,i}$ where $\beta$ is a hyperparameter that modulates the trade-off between specification adherence and the original task objective. Finally, we obtain the action distribution by applying softmax over the adjusted logits: $p_{t+k}=\text{softmax}(\tilde{\mathbf{z}}_{t+k})$